# Specific heat of Gd3+- and Eu2+-based magnetic compounds

### v2

@@ IV Effect of anisotropy
In this Section, we address the role of anisotropy first in the mean-field approximation and then incorporate the effect of quantum fluctuations.
@@ IV.1 Mean-field Approximation
In this subsection, we discuss the effect of the term $H_{A}$ on the specific heat calculated in the mean-field approximation, in which it is assumed that each site is subjected to a constant effective magnetic field resulting from the interaction with the remaining sites. Specifically, the interactions appearing in $H_{0}$ [see Eq. (1)] are approximated as
@@
We assume that all sites are equivalent and there is no external magnetic field. For simplicity, we restrict the analysis to the easy axis case $K<0$ for which the spins point in the direction of the anisotropy axis $z$. In the next subsection, both signs of $K$ and the effects of quantum fluctuations will be considered. For an antiferromagnetic order, we rotate 180 degrees the spins pointing in the $-z$ direction around a perpendicular axis so that for all sites $\langle S_{z}\rangle\geq 0$. This expectation value serves as the order parameter for the magnetic phase.
The effective magnetic field $B_{\text{eff}}$ is given by
$\tilde{B}=g\mu_{B}B_{\text{eff}}=I\langle S_{z}\rangle\text{,~{} }\langle S_{z}\rangle=\sum\limits_{M=-J}^{J}\frac{Me^{-\beta E_{M}}}{Z}\text{,~{}}Z=\sum\limits_{M=-J}^{J}e^{-\beta E_{M}},$ (5)
@@
$k_{B}T_{C}^{\text{MF}}=I\langle S^{2}_{z}\rangle\text{, where~{}~{}}\langle S^{2}_{z}\rangle=\sum\limits_{M=-J}^{J}\frac{M^{2}e^{-\beta E_{M}}}{Z}.$ (7)
@@
$C=T\frac{dS}{dT}\text{,~{}~{}~{} }S=\sum\limits_{M=-J}^{J}p_{M}\ln p_{M}\text{,~{}~{}~{} }p_{M}=\frac{e^{-\beta E_{M}}}{Z}$ (8)
Figure 3: (Color online) Specific heat over temperature as a function of temperature for several anisotropies in the mean-field approximation.
In Fig. 3 we show the resulting effect of the axial anisotropy on the specific heat keeping the value of $k_{B}T_{C}=1$ constant, as the unit of energy. For $K=0$, the value of the intensity for a given $T_{C}$ is $I=3k_{B}T_{C}/[S(S+1)]$, leading to $I=4/21\sim 0.190$ for $k_{B}T_{C}=1$.
For $K=-0.02$, we obtain $I=0.1531$ from Eq. (7). This is because the expectation value $\langle S^{2}_{z}\rangle$ that enters Eq. (7) is larger for $K<0$ and should be compensated with a smaller $I$. In addition, since the values of high $S_{z}$ projection $|M|$ are favored, the effective magnetic field is more efficient than for $K=0$ in splitting the lowest energy levels and therefore the entropy and specific heat decrease for $T<0.4T_{C}$ with respect to the isotropic case. This is compensated by an increase in $C/T$ for $0.5T_{C}<T<T_{C}$.
For $K=-0.04$, $I$ decreases further to $I=0.1263$, and the changes in the specific heat are more marked in the same direction. For large negative $K$, $I=1/(7/2)^{2}\sim 0.0816$.
@@ IV.2 Beyond Mean-field
In this subsection we discuss the effect of anisotropy in cases where the fluctuations of the effective field are also included.
Figure 4: (Color online) Specific heat over temperature as a function of temperature for several anisotropies for the simple cubic structure including 6 NN and 8 third NN with $J_{\delta}=-1$.
In Fig. 4 we show how the specific heat for a structure with $z=14$ is modified by the addition of $H_{A}$. The results are qualitatively the same as in the mean-field case for $T<T_{C}$. Addition of the anisotropy term shifts the entropy from the region $T<0.5T_{C}$ mainly to the region $0.5T_{C}<T<T_{C}$, and the effect is more pronounced for negative $K$.
However, in contrast to the mean-field case, addition of the anisotropy term decreases the specific heat for $T>T_{C}$. This reduction is expected because the specific heat in this region primarily results from quantum fluctuations, which are diminished by the energy level splitting induced by $H_{A}$.
@@ V Comparison of different structures with the same effective number of neighbors
Figure 5: (Color online) Specific heat over temperature as a function of temperature for several structures with different values of $r$ defined by Eq. (3). Top left: several structures with $J_{\delta}=-1$ and different $r$. The rest of the panels show the effect of changing $J_{\delta}$ keeping the same $r$.
In Section III, we discussed the effect of increasing the number of neighbors $z$ in the specific heat, while keeping $J_{\delta}=J_{1}$. Here we investigate the effect of changing $J_{\delta}$ for first, second, and possibly third nearest neighbors, keeping the same $r$ [see Eq. (3)], which corresponds to the inverse of the effective number of neighbors $z$. Some parameter sets may be highly unrealistic. However, the objective of this section is to demonstrate that entirely different parameter sets, which result in the same effective $z$, produce remarkably similar specific heat behaviors.
In Fig. 5 (b) we show the comparison for a moderately small $z$: a cubic BCC structure with 6 NN and 8 third NN. We compare the case of NN interaction $J_{1}=-1$ and NNN interaction $J_{2}=0$ with $J_{1}=-1/8$ and $J_{2}=-1$. In spite of the radical change in the magnitude of the interactions, the specific heats are similar.
In Fig. 5 (c), we show a similar comparison for an FCC structure with different combinations of $J_{1}$ and $J_{2}$. Again the results are similar, except for the case in which the dominant interaction is changed drastically from NN to NNN. When the dominant interaction is that of NNN, the specific heat is larger and the difference is of the order of that between different $z$ for $J_{1}=-1$ (Fig. 5 (a)). Finally, in Fig. 5 (d), we show a similar comparison for a larger $z$. In this case, the agreement between the two curves is excellent.
In conclusion, except for extreme cases, the magnitude of the specific heat at $T\sim T_{C}/2$ is mainly determined by the effective number of neighbors $z=1/r$. This result is significant, because it allows us to predict the expected specific heat taking into account the effective number of neighbors with a relative independence of the detailed magnetic structure.
@@ VI Fits of experimental data
In this Section, we present fits of the specific heat of six compounds: GdNi3Ga9, GdPbBi, GdCu2Ge2, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2. Previously, the specific heat of GdCoIn5 has been fitted [19].
The fitting parameters are the known $T_{C}$, which determines the magnitude of the interactions $J_{\delta}$ assumed equal, the number of neighbors $z$ and for the Eu compounds, the anisotropy $K$. The Gd compounds could be fitted with $K=0$. The assumption of $J_{\delta}$ independent of $\delta$ is justified from the results of the previous Section, which show that different $J_{\delta}$ with the same effective number of neighbors $1/r$ show very similar specific heats. The fact that $K$ is important only for the Eu compounds, might be expected for the $\sim 25\%$ larger ionic radius of Eu2+ compared to that of Gd3+ [28], rendering the former more sensitive to crystal-field effects.
The results are presented in Fig. 6. The Gd compounds are ordered with increasing effective number of neighbors $z$ in the fit. The parameters of the fits used for the different compounds are listed in Table 1. Also indicated is the number of NN in the reported structure, denoted as $z_{0}$. Notations like 1+1+2 in $z_{0}$ indicate that there are four atoms located at three nearly identical distances.
Compound $T_{N}$ (exp,[K]) $z_{0}$ $z$(fit) $J$[K] $K$[K]
GdNi3Ga9 20 3 4 1.52 0
GdCu2Ge2 12 4 8 0.38 0
GdPbBi 13 12 10 0.33 0
GdNiSi3 22 2+2+2 26 0.18 0
Eu2Pd3Sn 47 1+1+2 14 0.66 -0.26
EuPdSn2 12 2+2 14 0.55 -0.44
@@
Table 1: Parameters used in the fits: $T_{N}$ is the Neel temperature, $z_{0}$ is the number of nearest neighbors in the reported structure, $z$(fit) is the effective number of neighbors in the fit, $J$ is the intensity of the interactions, and $K$ is the anisotropy.
GdNi3Ga9 [29, 30] clearly corresponds to a low coordination number ($z=4$). This result is expected, since the honeycomb lattice described in Section III with antiferromagnetic interactions (three intralayer and one interlayer for each site) corresponds to the structure of the material and provides a very good fit.
@@
For the rest of the structures, ferromagnetic interactions are assumed. As stated in Section III, as long as there is no significant frustration, the difference between ferromagnetic and antiferromagnetic interactions for $z>6$ is only noticeable at very low temperatures, well below the temperature range of interest in this study. The key features of the specific heat curves we analyze, such as the shape of the peak and the high-temperature tail, are primarily determined by the effective coordination number and the single-ion anisotropy, and these features are largely insensitive to the sign of the exchange interaction in the absence of strong frustration.
GdCu2Ge2 has a tetragonal structure near to a BCC one (Fig. 4.1 of Ref. [31]). Assuming superexchange interaction through Cu, the “BCC” like character is even more plausible. Using only 8 NN interactions of a BCC structure, the resulting specific heat is very near the experimental one, although slightly below (above) the experimental one for $T<T_{C}$ ($T>T_{C}$), suggesting that $z$ might be a little bit larger. If only the four NN were considered, the system would be two-dimensional (with a broadened transition near $T_{C}$) and this is incompatible with the observed specific heat.
GdPdBi [32] is a half-Heusler structure with antiferromagnetic coupling. We obtain a very good fit assuming a layered honeycomb lattice as for GdNi3Ga9, but with a larger number of effective neighbors (10). The fitting is better than using an FCC structure. This effect might be caused by a small distortion of the lattice. Note that in this case, the effective number of neighbors ($z=10$) is slightly lower than the number of nearest neighbors ($z_{0}=12$). This is likely due to a slight frustration, which enhances the relative impact of fluctuations.
For GdNiSi3 [33], a simple cubic structure with $z=26$ including neighbors up to the third coordination shell provides a very good fit of the data. For a three-dimensional behavior, $z>16$ is necessary to connect different planes.
@@
This completes the Gd compounds that we have considered. In a previous work [19], the specific heat of GdCoIn5 has been fitted with essentially the same procedure. The compound has a tetragonal structure. The authors have extrapolated to the thermodynamic limit, the specific heat that corresponds to a simple cubic lattice $z=6$ NN with antiferromagnetic interactions, providing an excellent fit of the data.
Now turning to Eu compounds, the specific heat of Eu2Pd3Sn [34] is very well fitted using a simple cubic structure with 6 NN and 8 third NN ($z=14$) with $J_{1}=J_{2}$ and $K=-0.4J_{1}$. Small deviations exist at very low temperatures, where our numerical method is not fully reliable. The point group at the Eu sites is $C_{1h}$, compatible with the presence of $H_{A}$ as discussed in the introduction.
@@
Finally, the same simple cubic structure, but with $K=-0.8J_{1}$ provides a very good fit of the specific heat of EuPdSn2 [35]. For temperatures very near $T_{C}$, the theoretical data are a little above (below) the experimental ones for $T<T_{C}$ ($T>T_{C}$). The point group at the Eu sites is $C_{2v}$, which also allows for the inclusion of the term $H_{A}$. The discrepancies near $T_{C}$ ($T_{N}$) likely originate from factors not explicitly included in our model, such as more complex magnetic ordering (e.g., a collinear amplitude-modulated structure) resulting from modulated interactions and crystal-field effects, as discussed in Ref. 15.

### v3

@@ IV Effect of anisotropy
In this section, we address the role of anisotropy first in the mean-field approximation and then incorporate the effect of fluctuations.
@@ IV.1 Mean-field Approximation
In this subsection, we discuss the effect of the term $H_{A}$ on the specific heat calculated in the mean-field approximation. In this approximation it is assumed that each site is subjected to a constant effective magnetic field resulting from the interaction with the remaining sites. Specifically, the interactions appearing in $H_{0}$ [see Eq. (1)] are approximated as
@@
The mean-field approximation assumes that all sites are equivalent (long-range magnetic order), which may not hold near the ordering temperature. We assume that there is no external magnetic field. For simplicity, we restrict the analysis to the easy axis (uniaxial anisotropy or anisotropy along a single axis) case $K<0$, for which the spins point in the direction of the anisotropy axis $z$. In the next subsection, both signs of $K$ and the effects of spatial and quantum fluctuations will be considered.
For an antiferromagnetic order, we rotate 180 degrees the spins pointing in the z-z direction around a perpendicular axis so that for all sites $\langle S_{z}\rangle\geq 0$. This expectation value serves as the order parameter for the magnetic phase.
The effective magnetic field $B_{\mathrm{eff}}$ is given by
$Z$ $=\sum\limits_{M=-J}^{J}e^{-\beta E_{M}}\mathrm{,~}\langle S_{z}\rangle$ $=\sum\limits_{M=-J}^{J}\frac{Me^{-\beta E_{M}}}{Z}\mathrm{,~}$
$\tilde{B}$ $=g\mu_{B}B_{\mathrm{eff}}=I\langle S_{z}\rangle,$ (6)
@@
$k_{B}T_{C}^{\mathrm{MF}}=I\langle S^{2}_{z}\rangle\mathrm{,where~~}\langle S^{2}_{z}\rangle=\sum\limits_{M=-J}^{J}\frac{M^{2}e^{-\beta E_{M}}}{Z}.$ (8)
@@
$C=T\frac{dS}{dT}\mathrm{,~~~}S=\sum\limits_{M=-J}^{J}p_{M}\ln p_{M}\mathrm{,~~~}p_{M}=\frac{e^{-\beta E_{M}}}{Z}$ (9)
Figure 3: Specific heat $C/T$ versus $T/T_{C}$ for the mean-field approximation with varying axial anisotropy $K$ (expressed in units of $k_{B}T_{C}$).
In Fig. 3 we show the resulting effect of the uniaxial anisotropy on the specific heat keeping the value of $k_{B}T_{C}=1$ constant, as the unit of energy. For $K=0$, the value of the intensity for a given $T_{C}$ is $I=3k_{B}T_{C}/[S(S+1)]$, leading to $I=4/21\sim 0.190$ for $k_{B}T_{C}=1$.
For K=0.02K=-0.02, we obtain $I=0.1531$ from Eq. (8). This is because the expectation value $\langle S^{2}_{z}\rangle$ that enters Eq. (8) is larger for $K<0$ and should be compensated with a smaller $I$. In addition, since the values of high $S_{z}$ projection $|M|$ are favored, the effective magnetic field is more efficient than for $K=0$ in splitting the lowest energy levels and therefore the entropy and specific heat decrease for $T<0.4T_{C}$ with respect to the isotropic case. This is compensated by an increase in $C/T$ for $0.5T_{C}<T<T_{C}$.
For K=0.04K=-0.04, II decreases further to $I=0.1263$, and the changes in the specific heat are more marked in the same direction. For large negative $K$, $I=1/(7/2)^{2}\sim 0.0816$.
In the mean-field approximation, anisotropy shifts entropy from low to intermediate temperatures, with $T_{C}$ determined by the balance between exchange interactions and anisotropy. Above $T_{C}$ and when $K\neq 0$ the specific heat is different from zero reflecting the onsite energy fluctuations induced by the CEF term.
@@ IV.2 Beyond Mean-field
Figure 4: Specific heat $C/T$ versus $T/T_{C}$ for the EuPdSn2(16) structure with varying axial anisotropy $K$ ($J_{\delta}=-1$). The critical temperatures for K=0.6,0.3,0,0.3,0.6K=-0.6,-0.3,0,0.3,0.6 are approximately $78,74,68.5,71.5,73$ respectively. To compare with Figure 3 the ratio $K/{k_{B}T_{c}}$ for these cases are .0077-.0077, 0.0040-0.0040,0,$0.0042$ and $0.0082$ respectively.
In this subsection, we discuss the effect of anisotropy in cases where the fluctuations of the effective magnetic field are also included. Figure 4 shows the specific heat for the EuPdSn2(16) structure (described in section VI) with varying anisotropy $K$ for moderate values ($|K|\lesssim J_{\delta}$). For $K=0$, the system is isotropic, while negative $K$ values introduce axial anisotropy, favoring alignment along the $z$-axis. $K>0$ introduces a plane anisotropy favoring alignment in the $xy$-plane. Qualitatively, the results for $T<T_{C}$ mirror those of the mean-field approximation, with anisotropy shifting entropy from the region $T<0.5T_{C}$ to $0.5T_{C}<T<T_{C}$, and the effect is more pronounced for negative $K$.
An anisotropy-induced energy gap arises from the splitting of spin states due to the axial anisotropy term $H_{A}=K[3S_{z}^{2}-S(S+1)]$. For negative $K$, this term favors alignment along the $z$-axis, creating a gap between low-energy $|S_{z}|$ states with large $|S_{z}|$ values. This restricts spin fluctuations above $T_{C}$, reducing entropy and specific heat. For positive $K$, the anisotropy favors in-plane alignment, leading to a inverse distribution of low-energy states but with similar effects on the specific heat. The QMC simulations explicitly capture these quantum and spatial fluctuations, which mean-field theory cannot account for.
The Schottky-like shoulder observed in the specific heat is attributed to the internal magnetic field which creates a low-energy gap in the spin excitation spectrum. This feature is particularly sensitive to the axial anisotropy. For positive $K$, the low-energy states are more closely spaced, leading to a smaller effect compared to negative $K$ where low-energy states are more separated. In this last case the Schottky-like shoulder is more affected by the anisotropy.
The entropy removed above $T_{C}$ by the anisotropy term is recovered between the Schottky-like shoulder and $T_{C}$.
@@ V Comparison of Different Structures with the Same Effective Number of Neighbors
Figure 5: Specific heat $C/T$ versus $T/T_{C}$ for various lattice structures with different values of effective neighbor count $z$ (defined by Eq. 4), with $J_{1}=-1$. The critical temperatures $T_{C}$ and $z$ values are tabulated in Table 3. Top panel: Diamond lattice with varying $J_{2}$ interactions ($z=4,6,8,10,12,14,16$ from bottom to top). Inset: $z$ as a function of $J_{2}$. Middle panel: Structures with $z=8$ or $12$ (diamond, BCC and FCC). Bottom panel: Structures with $z=6$ or $18$ (diamond, SC, BCC and FCC).
Our central finding is that the specific heat of magnetic systems is primarily determined by the effective number of neighbors $z$, rather than the specific lattice structure or interaction pattern. In Section III, we analyzed the impact of varying the number of neighbors $z$ on the specific heat, while keeping $J_{\delta}=J_{1}$. Here, we extend this analysis by investigating how changes in $J_{\delta}$ affect the specific heat, while maintaining the effective number of neighbors $z$ constant [see Eq. (4)]. The parameter $z$ quantifies the relative strength of spatial and quantum fluctuations. Physically, $z$ represents the effective number of neighbors that meaningfully contribute to the exchange field experienced by each spin, accounting for both geometric coordination and relative interaction strengths.
To illustrate this principle, we use a diamond lattice with interactions $J_{1}=-1$ and varying $J_{2}$ as a guide. This allows us to explore cases where $z$ varies between $4$ (just nearest neighbors) and $16$ (nearest and next-nearest neighbors). In Figure 5(a), we show how the specific heat changes with $J_{2}$, with $z$ values ranging from $4$ to $16$. The specific heat curves closely follow the results shown in Figure 2, with even the $z=8$ and $z=12$ cases being nearly degenerate.
In the middle panel (b) and bottom panel (c) of Figure 5, we compare these diamond and non-uniform BCC and FCC results against lattices with uniform couplings (SC(6), BCC(8), FCC(12), FCC(18)). The specific heat curves for diamond and uniform coupling lattices with the same $z$ are strikingly similar. Also the diamond lattice with $J_{2}=-0.155$ ($z=8$) matches the BCC lattice with $J_{2}=-8.0$ ($z=8$), and the diamond lattice with $J_{2}=-0.333$ ($z=12$) matches the FCC lattice with $J_{2}=-4.0$ ($z=12$) . Despite the drastic changes in interaction magnitudes, the specific heat curves align closely, underscoring the dominance of $z$ over detailed interaction strengths. Even for the FCC lattice with $J_{2}=-4$, where curves do not overlap, the comparison remains good.
V.1 Getting information beyond mean field
The key insight of our work is that the effective number of neighbors $z$ (defined in Eq. 4) is the dominant parameter determining the specific heat, rather than the specific lattice structure or interaction pattern. This allows us to extract valuable information about the underlying exchange interactions that goes beyond what can be obtained from mean-field theory alone.
To illustrate this, consider the diamond lattice with varying $J_{2}$ (Fig. 5). For different values of $J_{2}$, we obtain different effective number of neighbors $z$. Crucially, when we compare diamond lattices with different $J_{2}$ values but the same $z$ to reference lattices with uniform couplings (SC, BCC, FCC), we find that the specific heat curves match almost perfectly. This demonstrates that the specific heat is insensitive to the detailed interaction pattern as long as $z$ is the same.
This observation provides a powerful tool for experimentalists. By comparing experimental specific heat data to our reference curves (available in Ref. 31), one can extract the effective number of neighbors $z$ and, from $z$, infer the relative strengths of different exchange interactions. This is a significant advance over mean-field analysis, which cannot distinguish between different interaction patterns that yield the same critical temperature.
Let’s examine how this works in practice using the examples from Fig. 5:
  1. 1.
    For the diamond lattice with $J_{2}=-0.080$ ($z=6.03$), the specific heat matches SC(6). Using Eq. 4 with $z_{1}=4$ and $z_{2}=12$ for the diamond lattice, we can solve for $J_{2}/J_{1}$:
    $1/z=\frac{z_{1}J_{1}^{2}+z_{2}J_{2}^{2}}{(z_{1}J_{1}+z_{2}J_{2})^{2}}=1/6.$
    With $J_{1}=-1$, this yields $J_{2}/J_{1}\approx 0.0787$, which is very close to the actual value of $0.080$. This demonstrates that the effective $z$ parameter can be used to extract the ratio of interaction strengths with high accuracy. Using the critical temperature as an energy scale, we can even determine the absolute values of $J_{1}$ and $J_{2}$.
  2. 2.
    Diamond lattices with $J_{2}=-0.155$ ($z=8$) and $J_{2}=-0.333$ ($z=12$) both match BCC(8) and FCC(12) as specific heat with $z=8$ and $12$ are very close. This indicates that these lattices require next-nearest-neighbor interactions for an accurate description. The analysis limits the possible parameter sets to $J_{2}/J_{1}\approx 0.155$ or $J_{2}/J_{1}\approx 0.333$, providing valuable constraints on the interaction strengths.
  3. 3.
    As in the previous case BCC($J_{2}=-0.40$) and BCC($J_{2}=-8.0$) lead to multiple possibilities: if $z=8$, then $J_{2}/J_{1}=0$ or $8$; if $z=12$, then $J_{2}/J_{1}=0.390$ or $2.276$.
  4. 4.
    For FCC($J_{2}=-4.0$), only two possibilities exist: $J_{2}/J_{1}=4$ ($z=12$) or $J_{2}/J_{1}=12.3$ ($z=8$).
  5. 5.
    For BCC($J_{1}=J_{2}=-1.0$, $J_{3}=-0.175$), the specific heat matches FCC(18) and lattices with $z=16$. This indicates that both nearest-neighbor (8) and next-nearest-neighbor (6) interactions are relevant. Considering interactions up to third order leads to a continuum of possibilities that can be further constrained by additional data.
The remarkable universality of specific heat curves across different lattice types with the same z suggests that the thermodynamics of these systems is primarily governed by the statistical properties of the exchange field distribution rather than detailed lattice geometry. As long as the exchange interactions are unfrustrated and the effective field distribution has the same variance-to-mean ratio (characterized by z), the thermodynamic behavior becomes universal.
This approach provides a systematic way to extract information about the underlying exchange interactions from specific heat measurements. It goes beyond mean-field theory by providing bounds for the number of relevant couplings and, in some cases, a limited set of options for interaction strengths. This is particularly valuable for experimentalists who often have limited information about the detailed magnetic interactions in their materials.
Lattice $J_{1}$ $J_{2}$ $J_{3}$ $z$ $T_{C}$ (K)
Diamond -1.0 -0.080 0.0 6.03 18.1
Diamond -1.0 -0.155 0.0 8 23.1
Diamond -1.0 -0.236 0.0 10 27.6
Diamond -1.0 -0.333 0.0 12 33.3
Diamond -1.0 -0.472 0.0 14 41.4
Diamond -1.0 -3.527 0.0 14 201.
BCC -1.0 -8.0 0.0 8 231.
BCC -1.0 -0.40 0.0 12 43.1
BCC -1.0 -1.0 -0.175 18 70.4
FCC -1.0 -4.0 0.0 12 151.
Table 3: Critical temperatures for Hamiltonian models with unequal interactions. All data corresponds to $N=20$.
The practical significance of this finding is substantial. By comparing experimental specific heat to our reference curves, one can extract the effective number of neighbors $z$ and, from $z$, infer the relative strengths of different exchange interactions. This information is inaccessible through mean-field analysis alone and can guide the development of more accurate theoretical models. For example, if a specific heat of a SC compound matches our reference curve for $z=8$ (and so it also matches $z=12$), we can immediately infer that the system likely has a combination of nearest-neighbor and next-nearest-neighbor interactions with a ratio $J_{2}/J_{1}\approx 0.08$ ($z=8$) or $J_{2}/J_{1}\approx 0.25$ ($z=12$).
This framework represents a significant advance in our ability to interpret magnetic specific heat data for 4f7 systems, as demonstrated in Section VI where we apply these principles to fit experimental data for six different compounds.
@@ VI Fits of experimental data
In this section, we present fits of the specific heat of six compounds: GdNi3Ga9, GdCu2Ge2, GdPbBi, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2 (see Figure 6). The fitting procedure uses the experimentally determined critical temperature $T_{N}$ as a fixed input, with the effective number of neighbors $z$ and axial anisotropy $K$ as free parameters. For Gd compounds, $K=0$ was used as the anisotropy term was found to be negligible. The assumption of uniform exchange interactions ($J_{\delta}=J$) is justified by our theoretical results showing that different $J_{\delta}$ values with the same effective number of neighbors $z$ yield nearly identical specific heat curves (see Section V).
The quality of the fits is excellent. For comparison, a mean-field approach would fail to reproduce the specific heat curves, particularly the shape of the peak and the high-temperature tail (see MF specific heat in Figures 2 and 3).
The parameters of the fits for the different compounds are listed in Table 4. The table also shows $z_{0}$. It represents the minimum number of nearest neighbors required to maintain 3D connectivity in the reported crystal structure. The effective number of neighbors $z$ accounts for both geometric coordination and relative interaction strengths, which can differ from the nominal $z_{0}$ due to longer-range interactions.
Compound $T_{N}$ (exp,[K]) $_{0} zz$(fit) |J||J|[K] KK[K]
GdNi3Ga9 20 4 4 1.5 0
GdCu2Ge2 12 12 12 0.20 0
GdPbBi 13 12 12 0.26 0
GdNiSi3 22 17 33 0.15 0
Eu2Pd3Sn 47 4 14 0.71 -0.28
EuPdSn2 12 8 22 0.11 -0.09
@@
Table 4: Parameters used in the fits: $T_{N}$ is the Neel temperature, $z_{0}$ is the number of nearest neighbors in the reported structure (compatible with 3D connectivity), $z$ (fit) is the effective number of neighbors, $J$ is the interaction strength, and $K$ is the anisotropy. This value correspond to $J_{1}=J_{2}$ and $J_{3}=0.27J_{1}$, see main text.
GdNi3Ga9 [33, 34] exhibits a layered honeycomb structure with $z_{0}=4$ (three intralayer and one interlayer nearest neighbors), consistent with $z=4$ found in the fit. The fit (Figure 6a) reproduces the experimental specific heat curve with excellent accuracy, with the peak at $T_{C}=20$ K matching the experimental value. The small $z$ value indicates significant spatial and quantum fluctuations, which is consistent with the observed broadening of the specific heat peak. As LHc(4,AF) lattice has $T_{N}/{|J_{1}|}\sim 13$ while experimentally $T_{N}=20K$ we can deduce $J\sim 20$$13\sim 1.5K$ corresponding to an antiferromagnetic interaction.
@@
For the rest of the structures, ferromagnetic interactions are assumed. As stated in Section III, as long as there is no significant frustration, the difference between ferromagnetic and antiferromagnetic interactions for $z>6$ is only noticeable at very low temperatures, well below the temperature range of interest in this study. In figure 5 we show the specific heat of a BCC lattice with AF interactions that compares well with the FM case. The key features of the specific heat curves, such as the shape of the peak and the high-temperature tail, are primarily determined by the effective number of neighbors and the single-ion anisotropy. These features are largely insensitive to the sign of the exchange interaction in the absence of strong frustration. For this reason we compare the experimental Neel temperature ($T_{N}$) with the theoretical Curie temperature ($T_{C}$).
GdCu2Ge2 has a tetragonal structure[35]. The Gd sublattice adopts a body-centered tetragonal structure. As this lattice is quite elongated ($c/a\gg 1$) if only the first four nearest-neighbor interactions were considered, the system would become two-dimensional with a broadened transition near $T_{N}$. The four next-nearest neighbors still lay in the same plane as the NN. To recover a tridimensional connectivity compatible with the experiments at least the eight following neighbors must be considered. This leaves as a minimum option to consider the first and third nearest neighbors ($z_{0}=12$) interactions. As explained before, this should be similar to consider other lattice with $z=12$ as the FCC lattice. In panel b) we show that indeed the specific heat can be described by any of these lattices. If, to simplify, we consider $J_{1}$ ($z_{1}=4$) equal to $J_{2}$ ($z_{2}=4$) we get the equation
$1/z=\frac{(z_{1}+z_{2})J_{1}^{2}+z_{3}J_{3}^{2}}{((z_{1}+z_{2})J_{1}+z_{3}J_{3})^{2}}.$
This assumption leads to three possibilities:
  • $J_{3}=0$ ($z=8$) which we can disregard has it implies a two dimensional behavior
  • $J_{3}/J_{1}=3.7$ ($z=12$) which we can disregard has it seems unphysical to have a larger strength associated to a more distant bond.
  • $J_{3}/J_{1}=0.27$ ($z=12$) appears to be the only sensitive solution.
Without the actual simulation for this combination of couplings, we can only estimate the coupling’s strength. To get the strength of the coupling we use the MF result for the critical temperature (see section IV.1)
$\frac{k_{B}T_{C}}{\frac{S(S+1)}{3}}=(z_{1}+z_{2})|J_{1}|+z_{3}|J_{3}|=(8+8\times 0.27)|J_{1}|$
from which we obtain $|J_{1}|=|J_{2}|\sim 0.20$K and $|J_{3}|\sim 0.05$K. From this results we can not determine the signs of the coupling. But this antiferromagnetic compound at least one of these coupling should be antiferromagnetic.
GdPdBi [36] is a half-Heusler structure. Gd sublattice in this compound forms a FCC structure. We obtain a very good fit with the LHc(10) lattice. Using $z=10$ leads to $J2/J_{1}\sim 6.1$ which seems unphysical. As FCC(12) ($T_{C}/{|J_{1}|}\sim 50$) also accounts for the experimental data we use that lattice to estimate the interaction strength. As in the previous compound we suppose that a NN interaction alone scenario is closer to the compound. Experimentally $T_{N}=13$K so we deduce $|J|\sim 13$K/50=0.26/50=0.26K. For this antiferromagnetic compound an antiferromagnetic coupling is expected.
The crystal structure of GdNiSi3 is an orthorhombic system. Gd atoms forms a bilayered arrangement of Gd. Only interaction at the 7th nearest neighbor connects bilayers to provide a tridimensional connectivity ($z_{0}=17$). A good description of the experimental data is obtain considering up to 10th nearest-neighbor ($z=25$) while the best fit is seen considering up to 13th NN ($z=33$). This suggests that long range interactions, well beyond first neighbors, are relevant in this compound. Given the complexity of these interactions we only consider the uniform $J_{\delta}$ case. The GdNiSi3(33) lattice has $T_{C}/{|J_{1}|}\sim 145$ while experimentally $T_{N}=22$K so we can deduce $|J|\sim 22$K/145=0.15/145=0.15K. The large amount of couplings needed (up to 13 nearest-neighbors) certainly come from a distribution of interactions. This distribution could be determined with first principle calculations that considers at least this coordination.
@@
This completes the Gd compounds that we have considered. In a previous work [32], the specific heat of GdCoIn5 has been fitted with essentially the same procedure. The compound has a tetragonal structure. The authors have extrapolated to the thermodynamic limit the specific heat that corresponds to a simple cubic lattice $z=6$ nearest-neighbor with antiferromagnetic interactions, providing an excellent fit of the data.
Now turning to Eu compounds, the importance of $K$ for the Eu compounds may be attributed to the $\sim 25\%$ larger ionic radius of Eu2+ compared to Gd3+ [37], which makes the former more susceptible to crystal-field effects.
The Eu sublattice of Eu2Pd3Sn can be though as a stacking of corrugated hexagonal layers in the ab plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=4$. This involves up to third order nearest-neighbor interactions. To properly compare the specific heat of Eu2Pd3Sn [38] with a Hamiltonian model we have to consider up to 6th nearest-neighbor in this lattice leading to a coordination $z=14$. The point group at the Eu sites is $C_{1h}$, which allows for the inclusion of the anisotropy term $H_{A}$, as discussed in section I.2. So an anisotropy term with $K/|J_{1}|=-0.4$ was included. With this anisotropy the Eu2Pd3Sn(14) lattice has $T_{C}/{|J_{0}|}\sim 66$. Experimentally $T_{N}=47$K so we can deduce, assuming uniform couplings, $|J|\sim 47$K/66=0.71/66=0.71K.
@@
Finally, the Eu sublattice of EuPdSn2 can be though as a stacking of rectangular layers in the ac plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=8$. This involves up to 3 nearest-neighbor interactions. To properly compare the specific heat of EuPdSn2 [29] with a Hamiltonian model we have to consider up to 7th nearest-neighbor in this lattice leading to a coordination $z=22$. The point group at the Eu sites is $C_{2v}$, which also allows for the inclusion of the term $H_{A}$. So an anisotropy term with $K/|J_{0}|=-0.9$ was included. With this anisotropy the EuPdSn2 lattice has $T_{C}/{|J_{0}|}\sim 112$. Experimentally $T_{N}=12$K from which we deduce $|J|\sim 12$K/112=0.11/112=0.11K. The discrepancies near $T_{N}$ likely originate from factors not explicitly included in our model, such as more complex magnetic ordering (e.g., a collinear amplitude-modulated structure) resulting from modulated interactions and crystal-field effects, as discussed in Ref. 14.